# General Hamiltonian description of nonreciprocal interactions

### v3

@@ VI Discussion and Outlook
The Hamiltonian framework we propose has the potential to expand the scope of research on nonreciprocal systems and enable novel applications. In experiments that simulate classical and quantum models [23, 68], our construction offers a new way to engineer nonreciprocal dynamics using reciprocal Hamiltonian systems subject to a constraint. Promising future directions include developing kinetic Monte Carlo schemes to capture non-steady states [38], canonically quantizing the Hamiltonian embedding, or using it to construct statistical mechanics and thermodynamic descriptions of nonreciprocal systems.
@@
Acknowledgments.—We thank L. Hahn, A. McRoberts, M. Sonner, Y. Zhang, and H. Zhao for related discussions. MB was funded by the European Union (ERC, QuSimCtrl, 101113633). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was in part supported by the Deutsche Forschungsgemeinschaft via Research Unit FOR 5522 (project-id 499180199) and the cluster of excellence ct.qmat (EXC 2147, Project No. 390858490).
@@ Appendix A Constrained Hamiltonian embedding of nonreciprocal systems
In this section, we first outline a general approach for constructing Hamiltonian embeddings. We then give the explicit form of the constrained Hamiltonian for various models with nonreciprocal equations of motion.
We begin by emphasizing that, in its simplest form, (non-)reciprocity is a property defined for pairwise interacting systems. Let us denote by ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})$ the force exerted by particle $i$ at position $\mathbf{r}_{i}$ on particle $j$ at position $\mathbf{r}_{j}$. We consider translation invariant systems where ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})={\bf F}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends only on the distance between the two particles involved. In addition, we only consider momentum-independent forces, and thus leave out of consideration Lorentz-like forces.
@@
We now assume that the interaction forces are conservative, i.e., they can be derived from an interaction potential:
@@
Reciprocal interactions share the same (mutual) interaction potential, and hence obey the relation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}_{j\rightarrow i}$ which, therefore, implies Newton’s third law (Eq. (12)); it is common to use the simplified notation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}$. By contrast, nonreciprocal interactions obey $U^{\text{NR}}_{i\rightarrow j}\neq U^{\text{NR}}_{j\rightarrow i}$. Since each interaction function $U^{\text{NR}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$, $U^{\text{NR}}_{j\rightarrow i}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends on the positions of both particles involved, one cannot write down an energy function such that the Hamiltonian equations of motion for the corresponding conjugate variables are nonreciprocal; this is because the resulting force acting on particle $i$ will inevitably contain the gradients of both interaction potentials which would make the net force reciprocal.
@@
The notation $\langle ij\rangle$ denotes a summation over pairs of particles $i$ and $j$ (each counted once). Here the kinetic energy is quadratic in the momenta $\mathbf{p}_{j}^{\mathbf{r}},\mathbf{p}_{j}^{\mathbf{x}}$, and hence the system is inertial and the EOM is second order in time. Note the apparent similarity of the kinetic energy above to that of a particle in a magnetic field; however, a notable difference is that in the Hamiltonian $H$ above the minimal coupling is between $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{x}_{j})$ [as opposed to $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{r}_{j})$], etc.

### v4

@@ VI Discussion and Outlook
The Hamiltonian framework we propose has the potential to expand the scope of research on nonreciprocal systems and enable novel applications. In experiments that simulate classical and quantum models [33, 54, 17, 84], our construction can offer a new way to engineer nonreciprocal dynamics using reciprocal Hamiltonian systems subject to a constraint. Promising future directions include developing kinetic Monte-Carlo schemes to capture non-steady states [47], canonically quantizing the Hamiltonian embedding, or using it to construct statistical mechanics and thermodynamic descriptions [90, 52] of nonreciprocal systems.
Methods
Equivalence between Constrained Glauber Monte-Carlo and Langevin dynamics
The constrained Glauber Monte-Carlo algorithm with transition rates defined using the Hamiltonian embedding and the Langevin equation with generic nonreciprocal interactions, share the same Fokker-Planck equation. This equation describes the configuration space density distribution $\rho(\vec{\theta},t)$, with $\vec{\theta}=(\theta_{i},\dots,\theta_{L^{2}})$. We now demonstrate this equivalence explicitly.
First, we sketch the derivation of the Fokker-Planck equation from the Langevin equation. Denoting the force acting on spin $\theta_{i}$ by $F_{i}(\vec{\theta})$, and using Itô formalism, the equation of motion is
$\theta_{i}(t+\Delta t)-\theta_{i}(t)=F_{i}(\vec{\theta})\,\Delta t+\sqrt{2T}\,\Delta W_{i}(t)\;,$ (11)
where $\Delta W_{i}(t)$ is a Wiener increment obeying $\langle\Delta W_{i}\rangle=0$ and $\langle\Delta W_{i}\,\Delta W_{j}\rangle=\delta_{ij}\Delta t$. Following the standard steps [1], we arrive at the Fokker-Planck equation for the configuration-space probability density $\rho(\vec{\theta},t)$:
$\partial_{t}\rho(\vec{\theta},t)=-\sum_{i}\partial_{\theta_{i}}\left[F_{i}(\vec{\theta})\,\rho(\vec{\theta},t)\right]+T\sum_{i}\partial_{\theta_{i}}^{2}\rho(\vec{\theta},t).$ (12)
In what follows, we independently derive Eq. (12) from the Glauber dynamics with transition probability rates
$w(\theta_{i}\to\theta_{i}^{\prime})$ $=$ $\frac{1}{2}\left(1-\tanh\frac{\Delta E(\theta,\theta^{\prime})}{2T}\right),$ (13)
$\Delta E(\theta,\theta^{\prime})$ $=$ $\frac{1}{2}\big(H(\theta^{\prime},\varphi)-H(\theta,\varphi^{\prime})\big)\big|_{\text{constr.($\theta,\varphi$)}}$ (14)
defined by eliminating the auxiliary degree of freedom from the Hamiltonian embedding by enforcing the constraint [1]. We note that $T$ here is the temperature of the bath; the nonreciprocal system of interest is never in a thermal state at any point during its evolution.
In each Glauber update step $n$, we pick a spin at random and update it according to $\theta_{j}\to\theta_{j}^{\prime}=\theta_{j}+\delta$. The probability of selecting spin $j$ is uniformly distributed on the square lattice of linear size $L$, and is hence given by $1/L^{2}$. At the same time, the update angle $\delta$ is drawn uniformly from the interval $\Delta,\Delta]$, with probability density $1/(2\Delta)$.
The discrete-time master equation governing these Glauber dynamics is then given by
$\rho(\vec{\theta},n+1)-\rho(\vec{\theta},n)=\frac{1}{2\Delta}\int_{-\Delta}^{\Delta}\mathrm{d}\delta\frac{1}{L^{2}}\sum_{i}\Big[$ (15)
$\rho(\theta_{i}+\delta,\{\vec{\theta}\}_{j\neq i},n)\;w(\theta_{i}{+}\delta\to\theta_{i}){-}\rho(\vec{\theta},n)\;w(\theta_{i}\to\theta_{i}{+}\delta)\Big].$
Treating $\Delta\ll 1$ as a small parameter, we can expand the integrand to leading order in $\delta$ using
$\rho(\theta_{i}{+}\delta,\{\vec{\theta}\}_{j\neq i},n)$ ${=}$ $\rho(\vec{\theta},n){+}\partial_{\theta_{i}}\rho(\vec{\theta},n)\delta{+}\frac{1}{2}\partial_{\theta_{i}}^{2}\rho(\vec{\theta},n)\delta^{2}$
${+}\mathcal{O}(\delta^{3})\;,$
$w(\theta_{i}{\to}\theta_{i}{+}\delta)$ ${=}$ $\frac{1}{2}{+}\frac{1}{4T}F_{i}(\vec{\theta})\delta{+}\frac{1}{8T}\partial_{\theta_{i}}F_{i}(\vec{\theta})\delta^{2}$
${+}\mathcal{O}(\delta^{3})\;,$
$w(\theta_{i}{+}\delta\to\theta_{i})$ ${=}$ $\frac{1}{2}{-}\frac{1}{4T}F_{i}(\vec{\theta})\delta{-}\frac{1}{8T}\partial_{\theta_{i}}F_{i}(\vec{\theta})\delta^{2}$ (16)
$+\mathcal{O}(\delta^{3})\;,$
where the force acting on spin $i$ is obtained from the embedding Hamiltonian under the constraint as
$F_{i}(\vec{\theta})$ $=$ $-\frac{1}{2}\left(\partial_{\theta_{i}}-\partial_{\varphi_{i}}\right)H(\vec{\theta},\vec{\varphi})\big|_{\text{constr.($\theta,\varphi$)}}$ (17)
$=$ $-\partial_{\theta_{i}}H(\vec{\theta},\vec{\varphi})\big|_{\text{constr.($\theta,\varphi$)}}\;.$
We emphasize that one first computes the derivatives and only then applies the constraint.
Noticing that linear terms in $\delta$ vanish since the integral limits are symmetric, we obtain to leading order in $\Delta$:
$\rho(\vec{\theta},n+1){-}\rho(\vec{\theta},n)\approx$ (18)
$\frac{1}{2\Delta L^{2}}\int_{-\Delta}^{\Delta}\delta^{2}\mathrm{d}\delta\sum_{i}\frac{1}{4T}\partial_{\theta_{i}}\Big[-\rho(\vec{\theta},n)F_{i}(\vec{\theta})+T\partial_{\theta_{i}}\rho(\vec{\theta},n)\Big]$
$=\frac{1}{L^{2}}\frac{\Delta^{2}}{12T}\sum_{i}\partial_{\theta_{i}}\Big[-\rho(\vec{\theta},n)F_{i}(\vec{\theta})+T\partial_{\theta_{i}}\rho(\vec{\theta},n)\Big]+\mathcal{O}(\Delta^{4}).$
Recalling that in Eq. (11) all $L^{2}$ spins get updated simultaneously at each time step, we define dimensionless Monte-Carlo time as
$t_{\textrm{MC}}=\frac{n}{L^{2}}.$ (19)
In the limit of large system sizes, $L^{2}\gg 1$, the corresponding time increment $\mathrm{d}t_{\mathrm{MC}}=1/L^{2}$ is small, and Monte-Carlo time becomes continuous; we obtain
$\rho(\vec{\theta},n+1)-\rho(\vec{\theta},n)=\partial_{t_{\mathrm{MC}}}\rho(\vec{\theta},t_{\mathrm{MC}})\;\mathrm{d}t_{\mathrm{MC}}.$ (20)
This allows us to formally define a continuous-time limit for the master equation:
$\frac{\partial\rho(\vec{\theta}{,}t_{\textrm{MC}})}{\partial t_{\textrm{MC}}}$ ${=}$ $\frac{\Delta^{2}}{12T}\sum_{i}\partial_{\theta_{i}}\Big[{-}\rho(\vec{\theta}{,}t_{\mathrm{MC}})F_{i}(\vec{\theta}){+}T\partial_{\theta_{i}}\rho(\vec{\theta}{,}t_{\mathrm{MC}})\Big]$ (21)
$+\mathcal{O}(\Delta^{4}).$
In the small-update regime $\Delta\ll 1$, Eq. (21) is equivalent to the Fokker-Planck equation derived earlier for the Langevin dynamics, cf. Eq. (12), except for the prefactor $\Delta^{2}/(12T)$.
In units where the friction coefficient is set to unity, it follows from the equations of motion that energy has the same unit as frequency (or inverse time) [see, e.g., Eq. (1)]. Since the only model-independent energy scale in the Glauber rate $w$ is the temperature $T$ (which also has units of energy), the physical time unit in Eq. (21) is set by the inverse bath temperature $T$. Hence, Monte-Carlo time $t_{\text{MC}}$ is related to the physical (Langevin) time $t_{\text{Lan}}$ by
$t_{\text{Lan}}=\frac{\Delta^{2}}{12T}t_{\text{MC}}.$ (22)
Therefore, starting with Glauber dynamics using the transition rates obtained from the Hamiltonian embedding, we derived the Fokker-Planck equation (12) which describes the Langevin dynamics of the original nonreciprocal system. This establishes the equivalence between the two approaches. We present a detailed formal derivation that uses the extended embedding variables in the SM [1].
Finally, we note that the derivation above does not make use of the details of a particular model. For as long as one can construct a Hamiltonian embedding, the latter can be used to define transition rates $w$ using Eq. (13), and the equivalence of the Langevin and Monte-Carlo descriptions follows.
Hamiltonian embedding for chase & run spin model
Consider XY spins on a square lattice of linear size $L$ with periodic boundary conditions. Denoting the lattice sites by i=(x,y)i{=}(x,y), the corresponding Langevin equation of motion for the chase-and-run spin model discussed in the main text reads as
$\dot{\theta}_{x,y}$ $=$ $-J_{\rightarrow}\sin(\theta_{x,y}-\theta_{x+1,y})-J_{\leftarrow}\sin(\theta_{x,y}-\theta_{x-1,y})$ (23)
$-J_{\text{rec}}[\sin(\theta_{x,y}-\theta_{x,y+1})+\sin(\theta_{x,y}-\theta_{x,y-1})]$
$+\sqrt{2T}\eta_{x,y}(t)\,.$
Here $J_{\leftarrow}{\neq}J_{\rightarrow}$ are the nonreciprocal interactions along the $x$-direction, and $J_{\text{rec}}$ denotes the reciprocal interactions along the $y$-direction (see main text). Temperature is denoted by $T$, and $\eta_{x,y}(t)$ adds white noise to the dynamics of each spin, cf. Eq. (1).
The spins are initialized in the perfect stripe state defined by
$\theta_{x,y}=\left\{\begin{array}[]{ll}0,&\text{for}\;x\in\left(\frac{L}{4},\frac{3L}{4}\right)\\ \frac{\pi}{2},&\mathrm{else.}\end{array}\right$.\, (24)
which is perturbed by the noise in the dynamics.
The Hamiltonian embedding for the nonreciprocal XY spins in Eq. (23) reads as
$H_{\text{tot}}$ $=$ $H_{SS}+H_{Sa}-H_{aa},$ (25)
$H_{SS}\!$ ${=}$ $\!\!{-}\!\!\sum_{x,y}(J_{\rightarrow}{+}J_{\leftarrow})[\cos(\theta_{x,y}{-}\theta_{x+1,y}){+}\cos(\theta_{x,y}{-}\theta_{x-1,y})]$
$\qquad+J_{\text{rec}}[\cos(\theta_{x,y}{-}\theta_{x,y+1}){+}\cos(\theta_{x,y}-\theta_{x,y-1})],$
$H_{Sa}\!$ ${=}$ $\!\!{-}\!\!\sum_{x,y}J_{\leftarrow}\cos(\theta_{x,y}{-}\varphi_{x+1,y}){+}J_{\rightarrow}\cos(\theta_{x,y}{-}\varphi_{x-1,y}),$
$H_{aa}\!$ ${=}$ $\!\!{-}\!\!\sum_{x,y}J_{\text{rec}}[\cos(\varphi_{x,y}{-}\varphi_{x,y+1}){+}\cos(\varphi_{x,y}{-}\varphi_{x,y-1})],$
where $\{\theta_{x,y},\varphi_{x^{\prime},y^{\prime}}\}=\delta_{x,x^{\prime}}\delta_{y,y^{\prime}}$ are conjugate variables.
A straightforward calculation shows that the Hamilton equations of motion derived from Eq. (25) reduce to the Langevin equation (23), when subject to the constraint
$\theta_{x,y}-\varphi_{x,y}=\pi\,.$ (26)
Thus, the energy difference used in the Glauber update rule can be evaluated following Eq. (6) as
$\Delta E(\theta_{x,y},\theta^{\prime}_{x,y})$ ${=}$ ${-}J_{\rightarrow}[\cos(\theta_{x,y}^{\prime}{-}\theta_{x{+}1,y}){-}\cos(\theta_{x,y}{-}\theta_{x{+}1,y})]$ (27)
${-}J_{\leftarrow}[\cos(\theta_{x,y}^{\prime}{-}\theta_{x-1,y}){-}\cos(\theta_{x,y}{-}\theta_{x-1,y})]$
${-}J_{\mathrm{rec}}\sum_{\sigma=\pm 1}[\cos(\theta_{x,y}^{\prime}{-}\theta_{x,y+\sigma})$
$\qquad\qquad\quad{-}\cos(\theta_{x,y}{-}\theta_{x,y+\sigma})]\;.$
We use this expression in the constrained Glauber simulations of the chase-and-run dynamics.
@@ Methods
Acknowledgments.—We thank L. Hahn, A. McRoberts, M. Sonner, Y. Zhang, and H. Zhao for related discussions. MB was funded by the European Union (ERC, QuSimCtrl, 101113633). YS sincerely thanks Hongzheng Zhao for hosting during the visit at Peking University. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work was in part supported by the Deutsche Forschungsgemeinschaft via Research Unit FOR 5522 (project-id 499180199) and the cluster of excellence ct.qmat (EXC 2147, Project No. 390858490).
@@ Appendix A Constrained Hamiltonian embedding of nonreciprocal systems
In this section, we first outline a general approach for constructing Hamiltonian embeddings, focusing on velocity-independent interaction forces. We then discuss the more general case of velocity-dependent interactions.
A.1 Velocity-independent interaction forces
We begin by emphasizing that, in its simplest form, (non-)reciprocity is a property defined for pairwise interacting systems. Let us denote by ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})$ the force exerted by particle $i$ at position $\mathbf{r}_{i}$ on particle $j$ at position $\mathbf{r}_{j}$. We consider translation invariant systems where ${\bf F}_{i\rightarrow j}(\mathbf{r}_{i},\mathbf{r}_{j})={\bf F}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends only on the distance between the two particles involved. In addition, in this first section, we only consider momentum-independent forces, and thus leave out of consideration Lorentz-like forces.
@@ A.1 Velocity-independent interaction forces
We now assume that the interaction forces can be derived from a one-way interaction potential:
@@
Reciprocal interactions are conservative and share the same (mutual) interaction potential, and hence obey the relation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}_{j\rightarrow i}$ which, therefore, implies Newton’s third law (Eq. (28)); it is common to use the simplified notation $U^{\text{R}}_{i\rightarrow j}=U^{\text{R}}$. By contrast, nonreciprocal interactions obey $U^{\text{NR}}_{i\rightarrow j}\neq U^{\text{NR}}_{j\rightarrow i}$, and are hence nonconservative. Since each interaction function $U^{\text{NR}}_{i\rightarrow j}(\mathbf{r}_{i}-\mathbf{r}_{j})$, $U^{\text{NR}}_{j\rightarrow i}(\mathbf{r}_{i}-\mathbf{r}_{j})$ depends on the positions of both particles involved, one cannot write down an energy function such that the Hamiltonian equations of motion for the corresponding conjugate variables are nonreciprocal; this is because the resulting force acting on particle $i$ will inevitably contain the gradients of both interaction potentials, which would make the net force reciprocal.
@@
The notation $\langle ij\rangle$ denotes a summation over pairs of particles $i$ and $j$ (each counted once). Here the kinetic energy is quadratic in the momenta $\mathbf{p}_{j}^{\mathbf{r}},\mathbf{p}_{j}^{\mathbf{x}}$, and hence the system is inertial and the EOM is second order in time. Note the apparent similarity of the kinetic energy above to that of a particle in a magnetic field; however, a notable difference is that, in the Hamiltonian $H$ above, the minimal coupling is between $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{x}_{j})$ [as opposed to $(\mathbf{p}_{j}^{\mathbf{r}},\mathbf{r}_{j})$], etc.